The power of a convex lens of focal length 5 cm is?

Grade:11

2 Answers

Arun
3 years ago
It is given that,
Focal length of the convex lens, f = 5 cm
To find,
The magnifying power of the lens.
Solution,
The formula of the magnifying power of a lens is given by :
D is the least distance of distinct vision, D = 25 cm
m = 1 + 25/5 = 1 + 5
m = 6
The magnifying power of the convex lens is 6. Hence, this is the required solution.
 
Vikas TU
3 years ago
Dear student 
Here, f=5cm,d=25cm,m=?
m=(1+d/f)=(1+25/5)=6.
///////////////////////////////////////////////